Hi, I can't seem to get the answer for this equation: -x-(13+4x)=-3(5-9x)+2
I have tried doing this and I end up getting x=13/18 but the answer is x=0. I know what I am doing wrong has something to do with the negative numbers on the left side of the equal sign but I don't know how or why you do it in that certain way.

Answer by KnightOwlTutor(293)

-x-(13+4x)=-3(5-9x)+2
remember (-)(-)=(+)
(-)(+)=(-)
Use the distributive property
-x-13-4x=-15+27x+2
simplify the equation and combine like terms
-5x-13=-13+27x
add 13 to both sides
add 5x to both sides
0=32x
The only way this equation can be correct is if x=0